# Decision-dependent Robust Charging Infrastructure Planning

### v1

@@ 2.2 Set, index, and parameters
  • $\Delta t$: Duration (mins) in each parking time interval
@@ 2.5 Objective function
The objective function consists of two stages: the charging planning and scheduling stages. The planning stage considers multiple types of charging infrastructure, such as slow and fast chargers. In addition to power capacity, we also model the SoC charging range for each type of charger, which is detailed in the following section. The operational part includes penalties for low battery SoC, which induces range anxiety and affects charging time during scheduling.
@@
where the low SoC penalty is calculated as the sum of the soft violation variable $v_{i,d,t}:=[0.3-b_{i,d,t}]^{+}$. This variable is greater than zero (i.e., $v_{i,d,t}>0$) if and only if the battery SoC level falls below 30 %. The charging time penalty is calculated as the sum of all the charging times across all zones and the time horizon to prevent low charger utilization.
@@ 2.6.1 Truck charging constraints
Constraint (1) ensures each truck uses at most one charger at any time. Constraint (2) enforces charging only during the scheduled parking window. Constraint (3) defines the binary indicator $\delta_{i,d,t}^{30}$ for the low SoC condition, where $\delta_{i,d,t}^{30}=0$ if and only if the battery level exceeds 30%. Here, $M_{1}$ is a large constant, and the effective range of $M_{1}$ is presented in Appendix 7. The parameter $\epsilon$ is a small positive number, and it ensures that the indicator equals one only when the battery SoC is strictly greater than 30%, thereby guaranteeing a unique waiting time when the battery SoC is exactly 30%.
@@ 2.6.3 Charging abandonment rules
At an industrial site, a truck abandons charging if its waiting time exceeds a maximum allowed waiting time $T^{\text{max}}_{i,d,t}$, as constraint (10). $M_{2}$ is a large constant to relax the waiting time constraint when $a_{i,d,t}$ equals to 1, in other words, when the trucks abandon charging. Notably, the maximum waiting time is a variable and it depends on the truck SoC due to the range anxiety and whether it is parked in special zones that allow for the overnight charging.
Based on experiences from the industrial partner, we formulate the following waiting and abandonment rules, considering range anxiety and overnight charging in the special zones. As constraint (11), if the SoC is greater than or equal to 30%, no waiting time is allowed as the driver has no range anxiety. If the SoC of a truck is below 30%, this will result in the driver’s range anxiety and would extend their waiting time to 0.5 hours before abandoning charging. Let the set $S^{Z}$ denote the parking time slots at any of the special zones for overnight charging. Constraint (12) sets the waiting time to be zero when the truck is outside the parking windows (i.e., no parking allowed) or in special zones (i.e., no waiting time limits). Constraint (13) regulates that the waiting time is always non-negative.
@@ 2.6.4 Battery dynamics constraints
Constraint (14) models EV battery dynamics considering charging power and energy consumption, and it should be noted that we do not use a binary variable to model battery discharging status, as it shall not happen during parking. Constraint (15) sets the initial SoC at $t_{0}$ and final SoC at time $t_{e}$ of the optimization horizon to be the same. Constraint (16) sets the maximum charging power limit considering the charging type $P^{j}$ and corresponding charging efficiency $\eta^{j}$. Constraint (17) regulates the charging time limit in each time window, which is greater than 5 mins and less than 30 minutes. Finally, constraint (18) restricts battery SoC within the maximum and minimum values.
@@ 2.6.5 Fast charger SoC restrictions
2.6.6 Decision-dependent uncertainty sets of parking duration
The effective charging duration $\hat{pp}$ is stochastic and influenced by the charging decision $y_{i,d,t}^{j}$, which is represented as binary variables. We constructed a box-shaped, decision-dependent uncertainty set for truck $i$ in zone $z$ at hour $t$. This is formulated as a robust linear constraint with affine decision dependence in terms of charging decisions, $y_{i,t,z}^{j}$.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot y_{i,d,t}^{j}\cdot\hat{pp}_{i,t,z},\quad\hat{pp}\in\Phi(y_{i,t,z}^{j})\quad\forall i,d,t\in\mathcal{S}$ (21)
The box-shaped uncertainty set $\Phi(y_{i,t,z}^{j})$ is defined as,
$\Phi=\left\{\hat{pp}\in\mathbb{R}^{i\times t}\ \Bigg|\ \frac{5}{30}\leq\hat{pp}\leq 1,\ \mu_{i,t,z}(y_{i,t,z}^{j})-\sigma_{i,t,z}(y_{i,t,z}^{j})\leq\hat{pp}\leq\mu_{i,t,z}(y_{i,t,z}^{j})+\sigma_{i,t,z}(y_{i,t,z}^{j})\right\}$ (22)
Where the bounds of the uncertainty set depend on the the minimum and maximum charging time, and calculated mean $\mu_{i,t,z}$ and standard deviation $\sigma_{i,t,z}$ associated with the charging decisions. The nominal moments are extracted from the historical data. Many charging trial data shows that fast charging could reduce the uncertainty of parking duration as the charging time becomes short. Therefore, we introduced a factor $\gamma$ to represent the reduction in uncertainty resulting from the fast charger usage.
$\mu_{i,t,z}(y_{i,t,z}^{j}):=\hat{\mu}_{i,t,z}\left(y_{i,t,z}^{0}+y_{i,t,z}^{1}\right)$ (23)
$\sigma_{i,t,z}(y_{i,t,z}^{j}):=\hat{\sigma}_{i,t,z}\left(\gamma^{0}\cdot y_{i,t,z}^{0}+\gamma^{1}\cdot y_{i,t,z}^{1}\right)$ (24)
Where $y^{0}$ and $y^{1}$ indicate charging decisions using slow or fast chargers, respectively. Notably, the decision-dependent RO is NP-complete, non-linear due to the multiplication of variables \@BBOPcite\@BAP\@BBNNohadani & Sharma (2018)\@BBCP. Here we proved that the proposed formulation is a special case that can be formulated as a robust linear constraint (33), and therefore the model can be formulated as a MILP problem. The proof is presented in Appendix 8.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}\cdot P^{j}\cdot\Delta t\cdot(\hat{\mu}_{i,t,z}\cdot y_{i,d,t}^{j}-\hat{\sigma}_{i,t,z}\cdot\gamma^{j}\cdot y_{i,t,z}^{j}),\quad\forall i,d,t\in\mathcal{S}$ (25)
@@ 2.6.7 Total number of chargers in each type and zone
3 Robust planning using the whole year dataset
@@
The planned EV infrastructure must remain robust across the year-round fleet operations. We compare two approaches to solving based on the whole-year dataset: day selection (DS) and iterative planning and scheduling (IPS).
@@
  • DS: We selected a certain number of representative days based on statistical criteria, such as the total fleet distance, the number of moving trucks, the total fleet energy consumption, and the total stopping time during a day. We also included days around the representative days to preserve the trip and battery cycle continuity.
@@
  • IPS: We first decomposed the full-year dataset into monthly subsets, allowing several overlapping days between months. The algorithm then plans charger deployment for the entire year by iteratively integrating scheduling and monthly planning. It starts with an initial charger allocation derived from the dataset and processes each month sequentially. If the existing infrastructure can satisfy the charging demand of this month, only charging scheduling is performed. Otherwise, the algorithm jointly determines the charging schedule and the additional chargers required, updates the infrastructure, and carries the new configuration forward to subsequent months.
4 Fix-and-optimize heuristic based on high charging demand forecasts
Directly solving the proposed planning and scheduling model can be computationally demanding with a long optimization horizon. In this paper, we propose a fix-and-optimize heuristic designated for the scheduling problem based on high charging demand forecasts. The fix-and-optimize heuristic has been applied to timetable \@BBOPcite\@BAP\@BBNFonseca et al. (2024)\@BBCP and job-shape scheduling \@BBOPcite\@BAP\@BBNLi et al. (2025)\@BBCP problems, where part of the solution is fixed while the remaining variables are free to form subproblems. In this paper, as presented in Fig. 4, we identified the parking intervals with high charging demand at an early stage based on the samples from the uncertainty set and pre-assigned fast-charging schedules, thereby accelerating the overall optimization process.
@@ 4 Fix-and-optimize heuristic based on high charging demand forecasts
Figure 4: The fix-and-optimize heuristic where fast-charging scheduling is pre-assigned and fixed based on the forecast high charging demands
As shown in Algorithm 1, we first initialize the master problem using the starting values $pp_{i,d,t}$ to solve the model. Next, we sample the stochastic parking duration $\hat{pp}_{i,d,t}$ from the uncertainty set $\Phi(\mu,\epsilon)$ to replace the initial values and compute the upper bounds of the charging power based on these samples. We then compare the required charging power against the updated upper bounds to identify any violations. A violation occurs if the required charging power exceeds the upper bound by more than a threshold $\kappa^{v}$ and the parking window does not belong to a special zone. The threshold $\kappa^{v}$ is initialized at two and decays exponentially across iterations. When a violation is detected, a fast-charging scheduling constraint $y_{i,1,d,t}=1$ is added for that parking window. These constraints are subsequently incorporated into the master problem, after which the model is reset and resolved from the first step.
Input: Parking duration samples $\hat{pp}_{i,d,t}$ drawn from the uncertainty set $\Phi(\mu,\epsilon)$
Output: Charger installation decisions $x_{z}^{j}$

### v2

@@ 2.2 Set, index and parameters
  • $\Delta t$: Duration (mins) in each parking time interval.
@@ 2.4 Uncertainty variables
  • $\bar{Y}_{c}^{j}\in\mathbb{R}^{+}$: The expected cluster-wise charging decision with the cluster $c$.
@@ 2.5 Objective function
The objective function includes the charging planning and scheduling. The charger planning accounts for multiple charging infrastructure types, including slow and fast chargers. In addition to power capacity, we also model the SoC charging range for each type of chargers, which are detailed in the following section. The scheduling part includes penalties for low battery SoC, inducing the range anxiety and the total charging time.
@@
where the low SoC penalty is calculated as the sum of the soft violation variable $v_{i,d,t}:=[0.3-b_{i,d,t}]^{+}$. This variable is greater than zero (i.e., $v_{i,d,t}>0$) if and only if the battery SoC level falls below 30 %. The charging-time penalty is calculated as the sum of charging times across all zones and the time horizon to prevent low charger utilization (i.e., low average charging power).
@@ 2.6.1 Truck charging constraints
Constraint (1) ensures each truck uses at most one charger at any time. Constraint (2) enforces charging only during the scheduled parking window. Constraint (3) defines the binary indicator $\delta_{i,d,t}^{30}$ for the low SoC condition, where $\delta_{i,d,t}^{30}=0$ if and only if the battery level exceeds 30%. Here, $M_{1}$ is a large constant, and the effective range of $M_{1}$ is presented in Appendix 6. The parameter $\epsilon$ is a small positive number that ensures the indicator equals 1 only when the battery SoC is strictly greater than 30%, thereby guaranteeing a unique waiting time when the battery SoC is exactly 30%.
@@ 2.6.3 Charging abandonment rules
At an industrial site, a truck abandons charging if its waiting time exceeds a maximum allowed waiting time $T^{\text{max}}_{i,d,t}$, as constraint (10). $M_{2}$ is a large constant to relax the waiting time constraint when $a_{i,d,t}$ equals to 1, in other words, when the trucks abandon charging. Notably, the maximum waiting time varies with the truck’s range anxiety and with whether it is parked in special zones that permit overnight charging.
Based on the industrial partner’s experience, we formulate the following waiting and abandonment rules, accounting for range anxiety and overnight charging in special zones as follows. As constraint (11), if the SoC is greater than or equal to 30%, no waiting time is allowed as the driver has no range anxiety. If the SoC of a truck is below 30%, this will result in the driver’s range anxiety and would extend their waiting time to half an hour before abandoning charging. Let the set $S^{Z}$ denote the parking time slots at any of the special zones for overnight charging. Constraint (12) sets the waiting time to be zero when the truck is outside the parking windows (i.e., no parking allowed) or in special zones (i.e., no waiting time limits). Constraint (13) regulates that the waiting time is always non-negative.
@@ 2.6.4 Battery dynamics constraints
Constraint (14) models EV battery dynamics considering charging power and energy consumption, and it should be noted that we do not use a binary variable to model battery discharging status, as it shall not happen during parking. Constraint (15) sets the initial SoC at $t_{0}$ and final SoC at time $t_{e}$ of the optimization horizon to be the same. Constraint (16) sets the maximum charging power limit considering the charging type $P^{j}$ and corresponding charging efficiency $\eta^{j}$. Constraint (17) regulates the charging time limit in each time window, which is greater than 5 mins and less than 30 mins. Finally, constraint (18) restricts battery SoC within the maximum and minimum values.
@@ 2.6.6 Total number of chargers in each type and zone
2.7 Decision-dependent robust formulation considering parking duration uncertainty
The stochastic effective parking duration, denoted by the random variable $\hat{pp}_{i,d,t}$, is unknown at the time of scheduling. We assume that this duration is endogenous: it depends on the charging decision $y_{i,d,t}^{j}$. Specifically, we model the realized parking duration $(\hat{pp})_{i,d,t}^{\prime}$ as an affine transformation of the intrinsic (base) random duration $\hat{pp}_{i,d,t}$.
Let $k^{j}\geq 1$ be a scaling constant associated with charger type $j$ (e.g., representing the reduction in charging time due to fast charging). The decision-dependent duration is defined as:
$(\hat{pp})_{i,d,t}^{\prime}:=\hat{pp}_{i,d,t}\cdot\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)$ (22)
Since we enforce single charger usage ($\sum y^{j}\leq 1$), the term in the parentheses acts as a scalar scaling factor.
2.7.1 Feature-driven cluster-wise uncertainty sets
Assuming that we have a total number of $N$ historical records of parking duration, each is indexed by $\omega\in\Omega$, $\Omega=\{1,2,...,N\}$, and records a truck’s parking duration $pp^{\omega}$ and its feature vector $v^{\omega}\in\mathcal{V}$, where $\mathcal{V}$ is the feature space. Each feature vector encodes information for each parking duration, including contextual and numerical variables such as half-hour of day, day of year, and zone. The fleet information describes the trucks’ operational purposes and is categorized into eight types, such as drill-and-blast or production activities. For instance, the feature vector of a parking duration $\omega$ can be represented as,
$\boldsymbol{v}^{\omega}=$ $\{1:(00:00),20:(\text{The 20th day of the year 2023}),$
$4:(\text{Zone 4}),3:(\text{Production team}),255:(\text{Truck ID 255})\}$
We found that parking duration is strongly correlated with fleet groups, half-hour intervals, and zones, but only slightly related to the day of the year. This motivates us to partition the parking duration into $c\in\mathbb{N}$ clusters to represent its stochasticity.
Given the calendar effect of parking events, we propose using hierarchical clustering by reordering the feature vectors for each truck. We compute the average parking duration for each half-hour zone combination, aggregated across all days of the year. These averaged parking durations under half-hour zone combinations serve as the feature vectors for individual trucks. Let the newly arranged feature vector be $(\boldsymbol{v}^{i})^{{}^{\prime}}$. Using these time–zone features alongside fleet information, we perform hierarchical clustering to group trucks, using Euclidean distance as the similarity measure.
$(\boldsymbol{v}^{i})^{{}^{\prime}}=$ $\{1:(00:00;\text{Zone 1}),2:(00:30;\text{Zone 1}),...,N_{t}\cdot N_{z}:(24:00;\text{Zone 8})\}$
We next model the distribution of uncertain parking duration for all trucks or truck groups and construct the corresponding feature-driven uncertainty set. Assuming that the parking duration of each cluster is identically and independently distributed, we could calculate the empirical mean $\mu_{c}$ and variance $\sigma_{c}^{2}=\frac{\sum_{\omega\in\Omega_{c}}(\hat{pp}^{\omega}-\mu_{c})^{2}}{N_{c}}$. Considering the inherent bounds of parking duration, we build a box-shaped uncertainty set as follows,
$\mathbb{P}:=\Bigg\{\hat{pp}\;\Big|\;\begin{array}[]{l}\mathbb{E}_{\mathbb{P}}[\hat{pp}]=\boldsymbol{\mu}_{pp},\\[2.0pt] \mathbb{E}_{\mathbb{P}}\!\left[(\hat{pp}-\boldsymbol{\mu}_{pp})(\hat{pp}-\boldsymbol{\mu}_{pp})^{\mathsf{T}}\right]=\boldsymbol{\Sigma}_{pp},\\[2.0pt] \lVert\hat{pp}\rVert_{\infty}\leq\Delta_{c}\end{array}\Bigg\}$ (23)
where $\boldsymbol{\mu}_{pp}=(\mu_{c}(\boldsymbol{v}))_{c\in\mathcal{C}}$, and the covariance matrix $\boldsymbol{\Sigma}_{pp}=\text{diag}(\sigma_{c}^{2}(\boldsymbol{v})_{c\in\mathcal{C}})$ are feature-driven. The values of $\Delta_{c}$ are based on $\boldsymbol{\mu}_{pp}$ and $\boldsymbol{\Sigma}_{pp}$, which control the conservativeness of our formulation.
2.7.2 Decision-dependent uncertainty sets and moment definitions
We could construct a decision-dependent uncertainty set $\mathbb{P}$ (23) based on the first and second moments. Let $\boldsymbol{\mu}_{c}$ and $\boldsymbol{\Sigma}_{c}$ denote the empirical mean vector and covariance matrix of the base parking duration for a cluster $c$, derived from historical data without charging interventions. We set $\Delta_{c}$ to be $\boldsymbol{\mu}_{c}\pm\boldsymbol{\sigma}_{c}$.
[…]
Example 1: Affine dependency on instantaneous decisions.
The moments are scaled directly by the specific decision for that parking slot. As $y_{i,d,t}^{j}$ is a “here-and-now” decision variable fixed prior to uncertainty realization, the adjusted moments are:
$\mu_{c}(y_{i,d,t})$ $=\hat{\mu}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)$ (24)
$\Sigma_{c}(y_{i,d,t})$ $=\hat{\Sigma}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{y_{i,d,t}^{j}}{k^{j}}\right)^{2}$ (25)
Example 2: Affine dependency on cluster-level expected decisions.
Here, the scaling factor depends on the aggregate behavior of the fleet within a cluster $c$. Let $N_{c}$ be the total number of parking slots in cluster $c$. We define the empirical expected value of charging decision $\bar{Y}_{c}$ as:
$\bar{Y}_{c}^{j}:=\frac{1}{N_{c}}\sum_{(i,d,t)\in c}y_{i,d,t}^{j}$ (26)
Note that $\bar{Y}_{c}$ is a linear function of decision variables, not a stochastic expectation.
Proposition 1 Given that the empirical expected value $\bar{Y}_{c}$ for a cluster of parking duration $c$, the moments for any instance in this cluster are scaled uniformly:
$\mu_{c}(\bar{Y}_{c}^{j})$ $=\hat{\mu}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)$ (27)
$\Sigma_{c}(\bar{Y}_{c}^{j})$ $=\hat{\Sigma}_{c}\left(1-\sum_{j\in\mathcal{J}}\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)^{2}$ (28)
Proof of Proposition 1 Please see 7.1 of the Supporting information.
2.7.3 RO-DDU constraint and reformulations
To ensure schedule feasibility, we enforce a RO-DDU constraint on the charging power limit:
[…]
Theorem 1. Example 1: Affine dependency on instantaneous decisions. Constraint (30) is the reformulation of the linear power limit constraint (29) under the robust uncertainty set with affine dependency on instantaneous decisions $\boldsymbol{y}^{j}$. This formulation preserves the linearity of binary charging decisions, keeping the problem tractable.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{1}{k^{j}}\right)\boldsymbol{y}^{j}-\left(1-\frac{1}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{y}^{j}\right]$ (30)
Example 2: Affine dependency on cluster-level expected decisions. After applying decision-dependent moments (27) and (28), constraint (31) is the reformulation of the linear power limit constraint (29) under the robust uncertainty set with affine dependency on cluster-level expected decisions $\bar{Y}_{c}$.
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)\boldsymbol{y}^{j}-\left(1-\frac{\bar{Y}_{c}^{j}}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{y}^{j}\right]$ (31)
This formulation is nonlinear because it includes the product of the binary variable $\boldsymbol{y}^{j}$ and the continuous variable $\bar{Y}_{c}$.
Proof of Theorem 1. Please see 7.2 of the Supporting information.
2.8 Cluster-based Linearization (Example 2)
While statistically rigorous, constraint (31) involves the product of the continuous variable $\bar{Y}_{c}$ (the sum of binary variables normalized by $N_{c}$) and the binary decision variable $y^{j}$, introducing bilinear terms. To solve this within a MILP framework, we introduce an auxiliary variable $Z_{i,d,t}^{j}=\bar{Y}_{c}^{j}\cdot y_{i,d,t}^{j}$. Since $\bar{Y}_{c}\in[0,1]$ and $y_{i,d,t}^{j}\in\{0,1\}$, this product can be exactly linearized using the McCormick envelopes \@BBOPcite\@BAP\@BBNMcCormick (1976)\@BBCP:
$Z_{i,d,t}^{j}\leq\bar{Y}_{c}^{j}$ (32)
$Z_{i,d,t}^{j}\leq y_{i,d,t}^{j}$ (33)
$Z_{i,d,t}^{j}\geq\bar{Y}_{c}^{j}-(1-y_{i,d,t}^{j})$ (34)
$Z_{i,d,t}^{j}\geq 0$ (35)
Substituting $\boldsymbol{Z}^{j}$ into (31) renders the constraint linear in the optimization variables, albeit at the cost of increasing the problem size. We obtain the constraint (36).
$p^{ch}_{i,d,t}\leq\sum_{j\in\mathcal{J}}\eta^{j}P^{j}\Delta t\left[\boldsymbol{\hat{\mu}}_{c}^{\intercal}\left(1-\frac{1}{k^{j}}\right)\boldsymbol{Z}^{j}-\left(1-\frac{1}{k^{j}}\right)\boldsymbol{\hat{\Sigma}}^{1/2}_{c}\boldsymbol{Z}^{j}\right]$ (36)
3 Solution Algorithm
3.1 Robust planning using the whole year dataset
@@
The planned EV infrastructure must remain robust across year-round fleet operations, that is, capable of meeting trip-related energy consumption, given the limited parking duration. We compare two planning results under different scheduling operations, based on the selective or full dataset:
@@
In terms of data usage, data can be categorized as day selection (DS) or full year (FY). For DS, we selected a subset of representative days based on statistical criteria, including total fleet distance, number of moving trucks, total fleet energy consumption, and total stopping time per day. We also include days around the representative days to preserve the trip and battery-cycle continuity. In contrast, the FY method uses the full-year dataset directly.
@@
For scheduling operations, we consider the hindsight (HS) and iterative planning and scheduling (IPS) approaches. The HS method assumes perfect foresight and performs one-shot scheduling. The IPS method is similar to rolling-horizon scheduling. The full-period dataset is first decomposed into monthly subsets, allowing several overlapping days between periods. The algorithm then plans charger deployment for the entire period by iteratively integrating scheduling and planning at the sub-period level. It starts with an initial charger allocation derived from the dataset and processes each sub-period sequentially. If the existing infrastructure can satisfy the charging demand of this sub-period, only charging scheduling is performed. Otherwise, the algorithm jointly determines the charging schedule and the additional chargers required, updates the infrastructure, and carries the new configuration forward to subsequent sub-periods.
@@
Figure 4: The fix-and-optimize heuristic where fast-charging scheduling is assigned and fixed based on the forecast high charging demands
The proposed planning and scheduling model is difficult to solve when the optimization horizon is long, as in hindsight scheduling. We propose a fix-and-optimize heuristic for the scheduling problem based on high-charging-demand forecasts. It has been applied to the timetable \@BBOPcite\@BAP\@BBNFonseca et al. (2024)\@BBCP and job-shape scheduling \@BBOPcite\@BAP\@BBNLi et al. (2025)\@BBCP, where part of the solution is fixed while the remaining variables are free to form subproblems. As shown in Fig. 4, we first identified parking intervals with high charging demand and pre-assigned fast-charging schedules, thereby accelerating the overall optimization process.
3.2 Fix-and-optimize heuristic based on high charging demand forecasts
Input: Master problem $\mathcal{M}$ with deterministic parking duration $pp_{i,d,t}$; Parking duration clusters $c(i,t,z)\in\mathcal{C}$
Output: BEV charger installation decisions $x_{z}^{j}$